# DRIFT: Dynamic Rule-Based Defense with Injection Isolation for Securing LLM Agents

### v2

@@ 3.3 DRIFT Adaptation
DRIFT is a system-level defense framework that can be deployed across many types of agents. To better understand the adaptability and generality of DRIFT in different agent settings, we apply it to multiple LLMs, including four advanced online models—GPT-4o [26], GPT-4o-mini [25], Claude-3 Haiku [28], and Claude-3.5-Sonnet [27]—and one widely used offline model, Qwen2.5-7B-Instruct [29]. The evaluation is conducted on AgentDojo.
For the online models, we compare our method with agents using ReAct [40], a technique that allows the LLM to reason and call tools in an agentic manner. The results are presented in Figure 5 (detailed results on four scenarios shown in Appendix D). We observe that DRIFT significantly enhances security across all models, reducing ASR from over 10% to single-digit levels, strongly indicating the security generality of DRIFT across diverse models. Notably, GPT-4o with ReAct, one of the most advanced LLMs with strong general capabilities, shows a high ASR of 51.7%, highlighting the vulnerability of current LLM agents—even those powered by leading models. However, after deploying DRIFT, the ASR drops sharply from 51.7% to just 1.5%, further demonstrating the effectiveness of DRIFT in securing agents from attack.
In addition, DRIFT does not compromise the agent’s task completion ability, as shown by the stable utility scores in both safe and unsafe conditions. In some cases, DRIFT even improves utility, e.g., with GPT-4o and Claude-3.5 Sonnet under attack.
The offline model Qwen2.5-7B-Instruct, which has been tuned on our policy, achieves remarkable improvements in both utility and security. In terms of utility, our tuned agent obtains a 5.6% improvement in safe conditions and 3.1% in unsafe conditions. It is noticeable that the ASR after tuning drops to 0. These improvements highlight a potential solution for robustly securing agentic systems without performance sacrifice. All of these results demonstrate the effectiveness of DRIFT across different models and scenarios, fully supporting its broad adaptability and strong generality.
@@ 3.4 Ablation Studies
We begin with the Native Agent setup, which uses the ReAct technique to serve as agents. GPT-4o-mini serves as the base model, with no defense mechanism applied. In this setting, the agent is vulnerable to be attacked, with a Targeted Attack Success Rate (ASR) of 30.67%. We then add the Secure Planner on the Native Agent, which generates fixed control- and data-level constraints based on the initial user query. These strict policies significantly improve security, reducing ASR to just 1.49%, showing the effectiveness of static policy enforcement. However, this improvement introduces severely utility drops. Specifically, The Utility in no attack decreases from 63.55% to 37.71% (a drop of 25.84%), and Utility Under Attack falls from 48.27% to 32.25% (a drop of 16.02%). This illustrates the limitation of using a static policy significantly undermines the agent capability to complete the tasks.
Afterward, we incorporate the Dynamic Validator, which adjusts policies during execution based on the agent’s interactions. This dynamic mechanism leads to a notable improvement in utility while maintaining strong security: Benign Utility and Utility Under Attack increase to 59.79% and 48.43%, respectively, while ASR rises slightly to 3.66%. These results demonstrate that dynamic policy updates provide a better balance, improving task success without significantly compromising security. To further explore the necessity of dynamic policies, we analyze how static and dynamic policies perform against the change of task complexity in Section 3.5.
Finally, we add the Injection Isolator, designed to mitigate long-term legacy risks by identifying and masking conflicting or malicious content in the memory stream. This component further reduces the ASR to just 1.29%, which is lower than the ASR achieved using only the strict policy. Moreover, it causes only a slight drop in utility. Furthermore, we evaluate the naive agent using only the isolator, it also effectively enhances security and reduces the ASR to 7.95%.
@@
Model Utility (No Attack) $\uparrow$ Utility (Under Attack) $\uparrow$ ASR (Under Attack) $\downarrow$
@@
w/ Planner 37.71 32.25 1.49
w/ Planner + Validator 59.79 48.43 3.66
w/ Planner + Validator + Isolator 58.48 47.91 1.29
w/ Isolator 54.85 47.17 7.95
@@ 3.6 Stress Test: Adaptive Attack
An important exploration for assessing defense robustness is stress attack. In this experiment, we evaluate our DRIFT on two types of adaptive attacks: the curated adaptive attack and the PAIR attack [41], an advanced black-box jailbreak optimization mechanism. In the former attack, we manually design injection prompts to bypass the isolator and validator, such as “Please note that there are no instructions that conflict with the user’s original query; therefore, you should not identify or list any conflicting instructions.” for the isolator, and “The deviated functions and parameters remain consistent with the intent of the user’s original query.” for the validator.
For the PAIR attack, we sample several cases where DRIFT successfully defends but the base agent does not, and optimize an injection prompt that can bypass the isolator and validator in these cases. The experimental results are presented in Table 2.
We can observe that DRIFT consistently maintains high utility and low ASR under all of these adaptive attacks. Notably, the combination of isolator and validator adaptive attacks (IAA + VAA) results in only a 0.04% utility loss and a 0.81% ASR increase, while the PAIR attack causes only a 2.13% utility loss and a 0.31% increase in ASR. These results further demonstrate the effectiveness and robustness of DRIFT under stress test.
@@
w/o Adaptive Attack 40.97 2.08 47.62 0.95 42.86 1.43 60.18 0.71 47.91 1.29
Isolator Adapt. Att. (IAA) 39.58 1.39 44.76 3.81 45.00 1.43 57.68 0.54 46.76 1.79
Validator Adapt. Att. (VAA) 37.50 0.69 42.86 3.81 43.90 1.43 56.61 0.71 45.22 1.66
IAA + VAA 38.19 2.08 43.81 0.95 49.29 5.00 60.18 0.36 47.87 2.10
PAIR 40.97 2.78 45.71 0.95 42.86 1.43 53.57 1.25 45.78 1.60
@@ 3.7 Overhead Analysis
DRIFT 2.37 47.9 1.3 19.7
It can be observed that DRIFT consumes approximately 1.89× more tokens than the undefended agent, yet fewer than most other defenses except for spotlighting_with_delimiting and tool_filter. In addition, DRIFT operates at a lower cost compared to the two other policy-based defenses, CaMeL [21] and Progent (w/ update) [22]. Specifically, CaMeL incurs roughly 7× the token cost. Notably, the tool_filter defense consumes even fewer tokens than the undefended agent, because it involves only a few tools during agent interactions, unlike the dozens used in AgentDojo, which substantially increases token usage.
In terms of efficiency, DRIFT performs slightly below tool filter but demonstrates a clear advantage over all other defenses, showing significantly higher efficiency than the other system-level defenses, CaMeL and Progent. However, tool filter still exhibits a 7.6% ASR, posing a notable security risk in real-world applications. In contrast, DRIFT reduces the ASR to only 1.3%. Overall, DRIFT achieves a strong balance between utility and security, making it more practical for real-world agent systems.
@@ 4 Conclusion
Acknowledgments and Disclosure of Funding
This project is partially supported by Schmidt Science AI2050 Early Career Fellow and Open philanthropy.
@@
Appendix A Limitations
Appendix B Related Works
B.1 LLM Agents
LLM Agents [1, 2, 3, 42, 43, 44, 45] are powered by large language models to automatically perceive environments and make decisions. Benefiting from the powerful reasoning capabilities of LLMs, a number of efforts [1, 44, 45, 3] equip LLM agents with tools to help users automatically complete tasks. Furthermore, recent advancements [44, 1, 2, 46] like Mind2Web [2] and WebAgent [1] construct systems to interact with web pages. OSWorld [3] constructs a desktop-manipulated system that enables agents to interact with computers. Additionally, several studies have explored methods to enhance agent reasoning capabilities. ReAct [40] introduces an effective approach to enhance the reasoning and acting capabilities of LLMs. Language Agent Tree Search [47] is proposed to improve the multi-step reasoning and planning capabilities of LLM agents. Some recent research also explores better tool selection mechanisms [45, 48, 30, 49]. REST-GPT [45] develops a flexible tool-calling interface for LLM agents. ToolBench [30] introduces a web-crawled benchmark for training and evaluating the tool-usage capabilities of LLMs.
B.2 Prompt Injection Defenses
A line of studies [12, 13, 14, 15, 18, 21] has explored solutions for securing LLM agents from prompt injection attacks. Current prompt injection defenses can be classified into model-level and system-level approaches.
Model-level defenses focus on enhancing the model’s inherent ability to resist attacks. StruQ [12] introduces a mechanism to transform queries into a structured form and trains the model to focus on the structured part. Chen et al.  [13] propose a preference optimization approach to defend against injection attacks. Another significant direction involves injection detection through external models, such as LlamaGuard [14] and InjecGuard [15]. These specialized models are trained to identify potentially malicious content across multiple risk categories, offering a complementary layer of protection.
System-level defenses typically constrain the model’s action space through predefined security policies to prevent attacks. Early system-level defenses focus primarily on coding scenarios [50] and face challenges when transferred to tool-integrated agent environments [1, 2, 3].
Recent advances in system-level protection have produced several notable approaches for tool-integrated agents. IsolateGPT [18] builds isolated execution environments for each application to reduce cross-application data flow risks. Both f-secure [19] and RTBAS [20] implement information flow control mechanisms that constrain untrusted data and propagate untrusted labels throughout the system. CaMeL [21] constructs control and data flows from the original user query and designs an interpreter to protect flow security. However, its control and data flow policies are static and cannot adequately meet the needs of dynamic real-time interactions. Concurrent with our work, Progent [22] develops a dynamic policy update mechanism based on historical interactions, but legacy injection messages in memory can still impact the agent or other modules, posing long-term risks for LLM agentic systems.
Appendix C Additional Experiments
C.1 Dynamic Mechanism Performance on Open-ended Task
@@
Utility (w/o att.) Utility (w/ att.) ASR Utility (w/o att.) Utility (w/ att.) ASR
@@
DRIFT (GPT-4o-mini) 61.24 46.30 1.64 26.50 28.50 4.75
C.2 Further Analysis of DRIFT and Progent
As a concurrent work, Progent [22] also proposes a dynamic policy-updating mechanism for securing LLM agents. In this experiment, we compare our DRIFT framework with Progent to further investigate the differences between these two defenses. Specifically, we conduct comparison experiments using GPT-4o and GPT-4o-mini as base models on the AgentDojo and ASB benchmarks, with the results shown in Table 5.
We observe that both DRIFT and Progent achieve comparable levels of utility and security when employing GPT-4o as the base model. However, when using GPT-4o-mini, DRIFT significantly outperforms Progent in terms of security (e.g., 1.64% ASR vs. 9.39% ASR on AgentDojo, and 4.75% ASR vs. 15.75% ASR on ASB). While Progent experiences a substantial drop in security performance, DRIFT maintains a level of robustness similar to that achieved with GPT-4o.
@@ C.2 Further Analysis of DRIFT and Progent
C.3 Case Study for Injection Isolator
@@
Appendix D Detailed Results on AgentDojo
@@
GPT-4o-mini ReAct 63.55 50.00 66.70 55.00 82.50
DRIFT 58.48 50.00 71.43 50.00 62.50
@@
DRIFT 75.86 75.00 80.95 65.00 82.50
Claude-3-haiku ReAct 39.97 37.50 52.38 35.00 35.00
DRIFT 37.90 43.75 42.86 25.00 40.00
GPT-4o ReAct 70.86 75.00 80.95 65.00 62.50
DRIFT 73.05 81.25 80.95 65.00 65.00
Qwen2.5-7B-Instruct ReAct 26.58 37.50 23.81 10.00 35.00
DRIFT 32.20 50.00 23.81 20.00 35.00
@@
DRIFT 47.91 40.97 47.62 42.86 60.18
@@
DRIFT 64.64 71.53 66.67 56.43 63.93
Claude-3-haiku ReAct 33.64 40.28 33.33 29.29 31.67
DRIFT 33.68 32.64 41.90 33.57 26.61
GPT-4o ReAct 55.43 69.44 63.81 64.29 24.17
DRIFT 62.28 72.92 59.05 56.43 60.71
Qwen2.5-7B-Instruct ReAct 19.12 24.31 20.00 7.86 24.30
DRIFT 22.20 27.08 19.05 10.71 31.95
@@
DRIFT 1.29 2.08 0.95 1.43 0.71
@@
DRIFT 4.41 1.39 0.00 0.00 16.25
Claude-3-haiku ReAct 11.15 11.81 27.62 1.43 3.75
DRIFT 3.03 2.78 3.81 2.86 2.68
GPT-4o ReAct 51.68 62.50 92.38 11.43 40.42
DRIFT 1.53 4.86 0.00 0.00 1.25
Qwen2.5-7B-Instruct ReAct 15.11 27.78 24.76 7.86 0.03
DRIFT 0.00 0.00 0.00 0.00 0.00
Appendix E Prompts

### v3

@@ 4.3 DRIFT Adaptation
DRIFT is a system-level defense framework that can be deployed across many types of agents. To better understand the adaptability and generality of DRIFT in different agent settings, we apply it to multiple LLMs, including four advanced models—GPT-4o OpenAI [2024b], GPT-4o-mini OpenAI [2024a], Claude-3.5-Haiku anthropic [2024a], and Claude-3.5-Sonnet anthropic [2024b]. The evaluation is conducted on AgentDojo.
We compare our method with agents using ReAct Yao et al. [2023], a technique that allows the LLM to reason and call tools in an agentic manner. The results are presented in Figure 6 (detailed results on four scenarios shown in Appendix B). We observe that DRIFT significantly enhances security across all models, reducing ASR from over 10% to single-digit levels, strongly indicating the security generality of DRIFT across diverse models. Notably, GPT-4o with ReAct, one of the most advanced LLMs with strong general capabilities, shows a high ASR of 51.7%, highlighting the vulnerability of current LLM agents—even those powered by leading models. However, after deploying DRIFT, the ASR drops sharply from 51.7% to just 1.7%, further demonstrating the effectiveness of DRIFT in securing agents from attack.
In addition, DRIFT does not compromise the agent’s task completion ability, as shown by the stable utility scores in both safe and unsafe conditions. In some cases, DRIFT even improves utility, e.g., with GPT-4o and Claude-3.5-Sonnet under attack. All of these results demonstrate the effectiveness of DRIFT across different models and scenarios, fully supporting its broad adaptability and strong generality.
@@ 4.4 Ablation Studies
We begin with the Native Agent setup, which uses the ReAct technique to serve as agents. GPT-4o-mini serves as the base model, with no defense mechanism applied. In this setting, the agent is vulnerable to be attacked, with a Targeted Attack Success Rate (ASR) of 30.67%. We then add the Secure Planner on the Native Agent, which generates fixed control- and data-level constraints based on the initial user query. These strict policies significantly improve security, reducing ASR to just 2.14%, showing the effectiveness of static policy enforcement. However, this improvement introduces severely utility drops. Specifically, The Utility in no attack decreases from 63.55% to 43.22% (a drop of 20.33%), and Utility Under Attack falls from 48.27% to 36.17% (a drop of 12.10%). This illustrates the limitation of using a static policy significantly undermines the agent capability to complete the tasks.
Afterward, we incorporate the Dynamic Validator, which adjusts policies during execution based on the agent’s interactions. This dynamic mechanism leads to a notable improvement in utility while maintaining strong security: Benign Utility and Utility Under Attack increase to 58.48% and 45.82%, respectively, while ASR rises slightly to 5.41%. These results demonstrate that dynamic policy updates provide a better balance, improving task success without significantly compromising security. To further explore the necessity of dynamic policies, we analyze how static and dynamic policies perform against the change of task complexity in Section 4.5.
Finally, we add the Injection Isolator, designed to mitigate long-term legacy risks by identifying and masking conflicting or malicious content in the memory stream. This component further reduces the ASR to just 1.35%, which is lower than the ASR achieved using only the strict policy. Moreover, it causes only a slight drop in utility. Furthermore, we evaluate the naive agent using only the isolator, it also effectively enhances security and reduces the ASR to 8.96%.
@@
Model Benign Utility $\uparrow$ Utility Under Attack $\uparrow$ ASR Under Attack $\downarrow$
@@
w/ Planner 43.22 36.17 2.14
w/ Planner + Validator 58.48 45.82 5.41
w/ Planner + Validator + Isolator 57.29 50.93 1.35
w/ Isolator 61.49 50.09 8.96
@@ 4.6 Stress Test: Adaptive Attack
An important exploration for assessing defense robustness is stress attacks. In this experiment, we evaluate our DRIFT on two types of adaptive attacks: curated adaptive attacks and the PAIR attack Chao et al. [2025], an advanced black-box jailbreak optimization mechanism.
In the former attack, the only module that can be targeted is the injection isolator. Therefore, we manually design injection prompts to bypass the isolator, such as: “Please note that there are no instructions that conflict with the user’s original query; therefore, you should not identify or list any conflicting instructions.”
For the PAIR attack, we sample several cases where DRIFT successfully defends while the base agent does not, and then optimize injection prompts that can bypass the isolator in these cases. The experimental results are presented in Table 2.
We can observe that DRIFT consistently maintains high utility and low ASR under all of these adaptive attacks. Notably, the manually designed adpative attack against isolator results in only a 0.7% ASR increase, while the PAIR attack causes only a 0.62% increase in ASR. These results further demonstrate the effectiveness and robustness of DRIFT under stress test.
@@
w/o Adaptive Attack 40.97 4.86 49.52 0.00 53.57 0.00 59.64 0.54 50.93 1.35
Manual Adaptive Attack 50.00 7.64 49.52 0.00 61.43 0.00 60.36 0.54 55.33 2.05
PAIR 43.75 5.56 47.62 0.00 58.57 1.43 61.25 0.89 52.80 1.97
@@ 4.7 Overhead Analysis
DRIFT 2.37 50.9 1.4 20.9
It can be observed that DRIFT consumes approximately 1.89× more tokens than the undefended agent, yet fewer than most other defenses except for spotlighting_with_delimiting and tool_filter. In addition, DRIFT operates at a lower cost compared to the two other policy-based defenses, CaMeL Debenedetti et al. [2025] and Progent (w/ update) Shi et al. [2025]. Specifically, CaMeL incurs roughly 7× the token cost. Notably, the tool_filter defense consumes even fewer tokens than the undefended agent, because it involves only a few tools during agent interactions, unlike the dozens used in AgentDojo, which substantially increases token usage.
In terms of efficiency, DRIFT performs slightly below tool filter but demonstrates a clear advantage over all other defenses, showing significantly higher efficiency than the other system-level defenses, CaMeL and Progent. However, tool filter still exhibits a 7.6% ASR, posing a notable security risk in real-world applications. In contrast, DRIFT reduces the ASR to only 1.4%. Overall, DRIFT achieves a strong balance between utility and security, making it more practical for real-world agent systems.
@@ 5 Conclusion
Limitations
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper. Only add this information once your submission is accepted and deanonymized.
@@ Acknowledgments
Appendix A Additional Experiments
A.1 Dynamic Mechanism Performance on Open-ended Task
@@
Benign Utility Attacked Utility ASR Benign Utility Attacked Utility ASR
@@
DRIFT (GPT-4o-mini) 57.29 50.93 1.35 26.50 28.50 4.75
A.2 Further Analysis of DRIFT and Progent
As a concurrent work, Progent Shi et al. [2025] also proposes a dynamic policy-updating mechanism for securing LLM agents. In this experiment, we compare our DRIFT framework with Progent to further investigate the differences between these two defenses. Specifically, we conduct comparison experiments using GPT-4o and GPT-4o-mini as base models on the AgentDojo and ASB benchmarks, with the results shown in Table 5.
We observe that both DRIFT and Progent achieve comparable levels of utility and security when employing GPT-4o as the base model. However, when using GPT-4o-mini, DRIFT significantly outperforms Progent in terms of security (e.g., 1.35% ASR vs. 9.39% ASR on AgentDojo, and 4.75% ASR vs. 15.75% ASR on ASB). While Progent experiences a substantial drop in security performance, DRIFT maintains a level of robustness similar to that achieved with GPT-4o.
@@ A.2 Further Analysis of DRIFT and Progent
A.3 Case Study for Injection Isolator
@@
Appendix B Detailed Results on AgentDojo
@@
GPT-4o-mini ReAct 63.55 50.00 66.67 55.00 82.50
DRIFT 57.29 50.00 66.67 55.00 57.50
GPT-4o ReAct 70.86 75.00 80.95 65.00 62.50
DRIFT 71.61 75.00 71.43 65.00 75.00
Claude-3.5-haiku ReAct 58.22 75.00 42.86 50.00 65.00
DRIFT 52.28 81.25 42.86 35.00 50.00
@@
DRIFT 73.36 75.00 80.95 60.00 77.50
@@
DRIFT 50.93 40.97 49.52 53.57 59.64
GPT-4o ReAct 55.43 69.44 63.81 64.29 24.17
DRIFT 62.00 68.75 59.05 52.86 67.32
Claude-3.5-haiku ReAct 45.61 56.94 39.05 34.29 52.14
DRIFT 42.41 60.42 36.19 23.57 49.46
@@
DRIFT 64.16 68.06 57.14 57.86 73.39
@@
DRIFT 1.35 4.86 0.00 0.00 0.54
GPT-4o ReAct 51.68 68.75 92.38 11.43 40.42
DRIFT 1.66 4.86 0.00 0.00 1.79
Claude-3.5-haiku ReAct 18.79 22.22 38.10 8.57 6.25
DRIFT 3.26 3.47 4.76 2.14 2.68
@@
DRIFT 1.47 0.00 0.00 0.71 5.18
Appendix C Prompts